Main Street Service bought four Michelin tires from a wholesaler for $343.00 and sold the tires at a markup of 30% of the selling price. (a) For how much were the tires sold? (b) What is the rate of markup based on cost?


1
Expert's answer
2021-01-13T19:23:58-0500

a)

"C=\$343.00, M=S\times\dfrac{30\%}{100\%}=0.3S"

"S=C+M"

Substitute


"S=343+0.3S""0.7S=343""S=\$490"



b)



"M=0.3S=0.3(\$490)=\$147""\dfrac{M}{C}\cdot100\%=\dfrac{147}{343}\cdot100\%=42.857\%"

The rate of markup based on cost is "42.857\%."